OpenStudy (anonymous):

How many electrons in the same atom can share the quantum number n = 2? A. 2 B.6 C.8 D.10

thomaster (thomaster):

\(\huge2*n^2\) just fill in

OpenStudy (anonymous):

Would it be 8?

OpenStudy (anonymous):

n=2 can have l=0 and l=1. l=0 has 1 orbital and l=1 has 3 orbitals. each of the orbital has 2 electrons

thomaster (thomaster):

yup, \(\large2*n^2=2*4=8\)

OpenStudy (anonymous):

Thank you

thomaster (thomaster):

but this only works when n alone is given

thomaster (thomaster):

when the angular quantum number (l) is given, it doesn't work anymore and you'll have to calculate it in another way

OpenStudy (anonymous):

So if when the angular quantum number (l) is given, how would you calculate that?